### v2

# Convergence in probability of numerical solutions of a highly non-linear delayed stochastic interest rate model

Abstract
We examine a delayed stochastic interest rate model with super-linearly growing coefficients and develop several new mathematical tools to establish properties of its true and truncated EM solutions. Moreover, we show that the true solution converges to the truncated EM solutions in probability as the step size tends to zero. Further, we support the convergence result with some illustrative numerical examples and justify the result via efficient use of the Monte Carlo method to value some financial quantities.
Keywords: Stochastic interest rate model, delay, truncated EM method, Monte Carlo method, bond, option contract.
Mathematics Subject Classification: 65C05, 65C30, 91G30, 91G60
1 Introduction
Stochastic modelling of interest rates is important for calibrating and valuing financial products such as option contracts. There exist variety of stochastic interest rate models proposed by some authors in the existing literature to describe the evolution of interest rates. One of the well-known stochastic interest rate models, also known as the CIR model was proposed by Cox, Ingersoll and Ross in [5]. This model follows the dynamics
$dx(t)=\alpha(\mu-x(t))dt+\sigma\sqrt{x(t)}dB(t).$ (1)
Here $x=(x(t),t\geq 0)$ denotes the interest rate with initial value $x(0)=x_{0}$, $\alpha,\mu,\sigma>0$ and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The CIR model is mean-reverting and due to its square root diffusion factor, it can also avoid possible negative rates.
We observe in SDE (1) that the volatility term $\sigma$ is assumed constant. However, as supported by empirical studies, volatility is not constant but exhibits empirical features widely known as volatility skews and smiles that are prevalent in option markets (see, e.g., [7, 6]). To provide adequate description of the evolution of volatility skews and smiles, one of the main schools of thought proposed based on empirical findings is to model volatility as a time-lagged dependent variable. For instance, Arriojas et al. in [1] proposed the delay Black-Scholes model where the drift depends on a time-lagged dependent variable and, also, the volatility term is a function of a time-lagged dependent variable. This model is described by the dynamics
$dx(t)=\mu x(t-a)x(t)dt+g(x(t-b))x(t)dB(t)$ (2)
on $t\in[0,T]$ with initial value $\xi(t)\in[-\tau,0]$, where $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\mu,a,b>0$, $\tau=\max(a,b)$, x(ta)x(t-a) and x(tb)x(t-b) denote delays in $x(t)$, $g:\mathbb{R}\rightarrow\mathbb{R}$ is a continuous function and $B=(B(t),t\geq 0)$ is a scalar Brownian motion. The authors showed that the delay Black-Scholes model maintains the no-arbitrage property and the completeness of the market with the correct volatility skews and smiles. Furthermore, Mao and Sabanis in [13] introduced the delay geometric Brownian motion described by the dynamics
$dx(t)=rx(t)dt+v(x(t-\tau))x(t)B(t)$ (3)
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau,r>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $B=(B(t),t\geq 0)$ is a scalar Brownian motion, the volatility function $v$ depends on $x(t-\tau)$ and $x(t-\tau)$ denotes delay in $x(t)$. The authors studied the quantitative properties of this model where $v$ is merely local Lipschitz continuous and bounded. The authors provided numerical evidence to justify that the system of type (3) is a rich alternative model for an asset price process in a complete market characterised by volatility skews and smiles. The reader may also consult (e.g., [16, 9]) for financial models with features of past dependency.
In recent years, more empirical studies have shown that the most successful continuous-time models for interest rates are those models that allow the volatility of interest changes to be highly sensitive to the level of the rates (see, e.g., [3, 14]). This motivated Wu et al. in [15] to extend SDE (1) to include super-linear diffusion term described by the dynamics
$dx(t)=\alpha(\mu-x(t))dt+\sigma x(t)^{\theta}dB(t)$ (4)
on $t\geq 0$, where $\theta>1$. One notable unique feature of SDE (4) is that the solution $x(t)$ is a highly sensitive mean-reverting process. The authors established the convergence in probability of the EM solutions to the true solution. They justified the convergence result within Monte Carlo simulations to value the expected payoff of a bond and a barrier option. To further capture volatility skews and smiles and high non-linearities in the rate, the authors in [4] extended the generalised Ait-Sahalia interest rate model to include a volatility as a function of a time-lagged dependent variable described by
$dx(t)=(\alpha_{-1}x(t)^{-1}-\alpha_{0}+\alpha_{1}x(t)-\alpha_{2}x(t)^{\gamma})dt+v(x(t-\tau))x(t)^{\theta}dB(t)$ (5)
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\alpha_{-1},\alpha_{0},\alpha_{1},\alpha_{2},\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\gamma,\theta>1$ and $B=(B(t),t\geq 0)$ is a scalar Brownian motion, $v$ is a function of $x(t-\tau)$ and $x(t-\tau)$ denotes delay in $x(t)$. Under a monotone condition and the assumption that $v$ is local Lipschitz continuous and bounded, the authors proved the strong convergence of the truncated EM solutions to the true solution of SDDE (5) and justified that the strong convergence result can be applied to value a bond and a barrier option.
Therefore, in other to account for high non-linearities in the rates as well as the evolution of volatility skews and smiles, we consider it necessary to reformulate SDE (4) as SDDE with super-linearly growing drift and diffusion coefficients described by the dynamics
$dx(t)=\alpha(\mu-x(t)^{\gamma})dt+\sigma x(t-\tau)^{r}x(t)^{\theta}dB(t),$ (6)
on $t\geq 0$ with the initial value $\xi(t)$ on $t\in[-\tau,0]$, where $\tau>0$, $\xi:\Omega\rightarrow C([-\tau,0]:\mathbb{R})$, $\theta,r>0$ and $\gamma>1$. We observe that both the drift factor $x^{\gamma}$ and the diffusion factor $x^{\theta}y^{r}$ of SDDE (6) are growing super-linearly and thus violate the global Lipschitz and linear growth conditions. This is further complicated by the presence of the unbounded time-lagged dependent variable $y$. Therefore, it can be very challenging to obtain the solution of SDDE (6) by an analytical closed-form formula. To the best of our knowledge, there exist no relevant literature for numerical treatment of SDDE (6) either in the strong sense or weak sense. In this case, we recognise the need to examine the feasibility of the system of SDDE (6) from a viewpoint of financial applications. Therefore, we need an efficient numerical method with fast computational performance to estimate the solution. However, in most real-world applications, the explicit EM method is preferred to the implicit type due to its simple algebraic structure, cheap computational cost and acceptable convergence rate. It is well-known in [8] that the explicit EM scheme diverges in the strong mean-square sense at finite point for SDEs with super-linearly growing coefficient terms. In this work, we aim to construct a variant of the truncated EM method developed in [11] to estimate the true solution of SDDE (6) and show that the truncated EM solutions converge to the true solution in probability when the step size is sufficiently small. The remainder of the paper is organised as follows: We explore mathematical notations in Section 2. In Section 3, we study properties of the true solution of SDDE (6). We construct the truncated EM techniques to approximate SDDE (6) and study properties of the truncated EM solutions in Section 4. In Section 5, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We also provide illustrative numerical examples to support the convergence result and justify the result via efficient use of the Monte Carlo method to value a bond and a lookback put option in this section .
2 Mathematical preliminaries
Throughout this paper, unless specified otherwise, we employ the following notation. Let $\{\Omega,\mathcal{F},\mathbb{P}\}$ be a complete probability space with filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ satisfying the usual conditions (i.e., it is increasing and right continuous while $\mathcal{F}_{0}$ contains all $\mathbb{P}$ null sets), and let $\mathbb{E}$ denote the expectation corresponding to $\mathbb{P}$. Let $B=(B(t),t\geq 0)$, be a scalar Brownian motion defined on the above probability space. If $a,b$ are real numbers, then $a\vee b$ denotes the maximum of $a$ and $b$, and $a\wedge b$ denotes the minimum of $a$ and $b$. Let $\mathbb{R}=(-\infty,\infty)$ and $\mathbb{R}_{+}=(0,\infty)$. If $x\in\mathbb{R}$, then $|x|$ is the Euclidean norm. For $\tau>0$, let $C([-\tau,0];\mathbb{R}_{+})$ denote the space of all continuous functions $\xi:[-\tau,0]\rightarrow\mathbb{R}_{+}$ with the norm $\|\xi\|=\sup_{-\tau\leq t\leq 0}\xi(t)$. For an empty set $\emptyset$, we set inf $\text${inf $\emptyset$$\infty$. For a set $A$, we denote its indication function by $1_{A}$. Let the following scalar dynamics
$dx(t)=f(x(t))dt+g(x(t),x(t-\tau))dB(t)$ (7)
with initial value $x(t)=\xi(t)\in C([-\tau,0]:\mathbb{R}_{+})$ denote equation of SDDE (6) such that $f(x)=\alpha(\mu-x^{\gamma})$ and $g(x,y)=\sigma x^{\theta}y^{r}$, for all $x,y\in\mathbb{R}_{+}$. Let $C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$ be the family of all real-valued functions $V(x,t)$ defined on $\mathbb{R}\times\mathbb{R}_{+}$ such that $V(x,t)$ is twice continuously differentiable in $x$ and once in $t$. For each $V\in C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$, define the operator $LV:\mathbb{R}\times\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$ by
$LV(x,y,t)=V_{t}(x,t)+V_{x}(x,t)f(x)+\frac{1}{2}V_{xx}(x,t)g(x,y)^{2}$ (8)
for SDDE (6) associated with the function $V$, where $V_{t}(x,t)$ and $V_{x}(x,t)$ are first-order partial derivatives with respect to $t$ and $x$ respectively, and $V_{xx}(x,t)$, a second-order partial derivative with respect to $x$. With the operator $LV$ defined, then the Itô formula yields
$dV(x(t),t)=LV(x(t),x(t-\tau),t)dt+V_{x}(x(t),t)g(x(t),x(t-\tau))dB(t)$ (9)
almost surely. We should emphasise that $LV$ is defined on $\mathbb{R}\times\mathbb{R}\times\mathbb{R}_{+}$ while $V$ is defined on $\mathbb{R}\times\mathbb{R}_{+}$. Moreover, we impose the following standing hypotheses.
Assumption 2.1.
The parameters of SDDE (6) satisfy
$1+\gamma>2(r+\theta),$ (10)
where $\theta,r>0$ and $\gamma>1$.
Assumption 2.2.
There exist constants $D>0$ and $\ell\in(0,1]$ such that for all $\tau\leq s\leq t\leq 0$, the initial function $\xi$ satisfies
$|\xi(t)-\xi(s)|\leq D|t-s|^{\ell}.$ (11)
We introduce the following lemma for later use.
Lemma 2.3.
Let Assumption 2.1 hold. For any $R>0$, there exists a constant $G_{R}>0$ such that the coefficients of the SDDE (6) satisfy
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}\big(|x-\bar{x}|+|y-\bar{y}|\big)$ (12)
for all $x,y,\bar{x},\bar{y}\in\mathbb{R}$ with $\max\{|x|,|\bar{x}|,|y|,|\bar{y}|\}\leq R$.
Proof.
We only prove the result for $x\leq\bar{x}$ and $y\leq\bar{y}$; the general case follows similarly. First observe that
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))$
$=\alpha(\mu-x^{\gamma})-\alpha(\mu-\bar{x}^{\gamma})+\sigma x^{\theta}y^{r}-\sigma\bar{x}^{\theta}\bar{y}^{r}$
$=-\alpha(x^{\gamma}-\bar{x}^{\gamma})+\sigma(x^{\theta}y^{r}-\bar{x}^{\theta}\bar{y}^{r}).$
Using Young inequality, we have
$x^{\theta}y^{r}\leq x^{2\theta}+y^{2r},\quad\bar{x}^{\theta}\bar{y}^{r}\leq\bar{x}^{2\theta}+\bar{y}^{2r}.$
Hence,
$x^{\theta}y^{r}-\bar{x}^{\theta}\bar{y}^{r}\leq(x^{2\theta}-\bar{x}^{2\theta})+(y^{2r}-\bar{y}^{2r}).$
It follows that
$(f(x)-f(\bar{x}))+(g(x,y)-g(\bar{x},\bar{y}))$
$\leq-\alpha(x^{\gamma}-\bar{x}^{\gamma})+\sigma(x^{2\theta}-\bar{x}^{2\theta})+\sigma(y^{2r}-\bar{y}^{2r}).$
Applying the mean value theorem to each term yields
$|x^{\gamma}-\bar{x}^{\gamma}|$ $\leq\gamma\big(|x|^{\gamma-1}+|\bar{x}|^{\gamma-1}\big)|x-\bar{x}|,$
$|x^{2\theta}-\bar{x}^{2\theta}|$ $\leq 2\theta\big(|x|^{2\theta-1}+|\bar{x}|^{2\theta-1}\big)|x-\bar{x}|,$
$|y^{2r}-\bar{y}^{2r}|$ $\leq 2r\big(|y|^{2r-1}+|\bar{y}|^{2r-1}\big)|y-\bar{y}|.$
Therefore,
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|$
$\leq\Big[\alpha\gamma(|x|^{\gamma-1}+|\bar{x}|^{\gamma-1})+2\theta\sigma(|x|^{2\theta-1}+|\bar{x}|^{2\theta-1})\Big]|x-\bar{x}|$
$\quad+2r\sigma(|y|^{2r-1}+|\bar{y}|^{2r-1})|y-\bar{y}|.$
Since $|x|,|\bar{x}|,|y|,|\bar{y}|\leq R$, all coefficients are bounded. Hence there exists a constant $G_{R}>0$ such that
$|f(x)-f(\bar{x})|+|g(x,y)-g(\bar{x},\bar{y})|\leq G_{R}(|x-\bar{x}|+|y-\bar{y}|).$
This completes the proof. ∎
3 Properties of true solution
In this section, we study properties of the true solution to SDDE (6). Since SDDE (6) is a financial model, it is a natural requirement to show that the solution is always positive.
3.1 Existence of positive solution
The following theorem shows that the solution of SDDE (6) is positive almost surely.
Theorem 3.1.
Let Assumption 2.1 hold. Then for any given initial value
$\{x(t):-\tau\leq t\leq 0\}=\xi(t)\in C([-\tau,0]:\mathbb{R}_{+}),$ (13)
there exists a unique solution $x(t)$ to SDDE (6) and $x(t)>0$ almost surely.
Proof.
Since the coefficients of SDDE (6) satisfy local Lipschitz condition in $\tau,\infty)$, one can show by the standard truncation method that there exists a unique maximal local solution $x(t)$ on $\tau,\eta_{e})$ for any given initial value (13), where $\eta_{e}$ is the explosion time (see [12]). Let $k_{0}>0$ be sufficiently large such that
$\frac{1}{k_{0}}<\underset{-\tau\leq t\leq 0}{\min}|\xi(t)|\leq\underset{-\tau\leq t\leq 0}{\max}|\xi(t)|<k_{0}.$
For each integer $k\geq k_{0}$, we define the stopping time by
$\eta_{k}=\inf\{t\in[0,\eta_{e}):x(t)\not\in[1/k,k]\}.$ (14)
We observe that $\eta_{k}$ is increasing as $k\rightarrow\infty$. We set $\eta_{\infty}=\underset{k\rightarrow\infty}{\lim}\eta_{k}$, whence $\eta_{\infty}\leq\eta_{e}$ almost surely. In other words, we need to show that $\eta_{\infty}=\infty$ almost surely to complete the proof. We define a $C^{2}$-function $V:\mathbb{R_{+}}\rightarrow\mathbb{R_{+}}$ by
$V(x)=x^{\beta}-1-\beta\log(x),$ (15)
where $\beta\in(0,1)$. By applying the operator defined in (8) to (15), we compute
$LV(x,y)$ $=\beta(x^{\beta-1}-x^{-1})\alpha(\mu-x^{\gamma})+\frac{\sigma^{2}}{2}(\beta(\beta-1)x^{\beta-2}+\beta x^{-2})x^{2\theta}y^{2r}$
$=\alpha\mu\beta x^{\beta-1}-\alpha\mu\beta x^{-1}-\alpha\beta x^{\gamma+\beta-1}+\alpha\beta x^{\gamma-1}$
$+\frac{\sigma^{2}}{2}\beta(\beta-1)x^{2\theta+\beta-2}y^{2r}+\frac{\sigma^{2}}{2}\beta x^{2\theta-2}y^{2r}.$
We note that for $\beta\in(0,1)$, $\gamma+\beta-1>2\theta+2r+\beta-2\Rightarrow\gamma+1>2(r+\theta)$. So by Assumption 2.1 and for $\beta\in(0,1)$, $\alpha\beta x^{\gamma+\beta-1}$ leads and tends to -$\infty$ for large $x$. However, for small $x$, $\alpha\mu\beta x^{-1}$ leads and also tends to -$\infty$. Hence, we can find a constant $K_{0}$ such that
$LV(x,y)\leq K_{0}.$ (16)
For any $t_{1}\in[0,\tau]$, the Itô formula yields
$\mathbb{E}(V(x(\eta_{k}\wedge t_{1})))$ $\leq V(\xi(0))+\mathbb{E}\int_{0}^{\eta_{k}\wedge t_{1}}K_{0}ds$
$\leq V(\xi(0))+K_{0}\tau,$
for all $k\geq k_{0}$. Noting that
$\mathbb{P}(\eta_{k}\leq\tau)\leq\frac{\mathbb{E}[V(x(\eta_{k}\wedge t_{1}))]}{V(k)\wedge V(1/k)},$
we have
$\mathbb{P}(\eta_{k}\leq\tau)\leq\frac{V(\xi(0))+K_{0}\tau}{V(k)\wedge V(1/k)}.$
As $k\rightarrow\infty$, $\mathbb{P}(\eta_{k}\leq\tau)\rightarrow 0$ and hence, $\eta_{\infty}>\tau$ almost surely. For $t_{1}\in[0,2\tau]$, the Itô formula gives us
$\mathbb{E}(V(x(\eta_{k}\wedge t_{1})))$ $\leq V(\xi(0))+\mathbb{E}\int_{0}^{\eta_{k}\wedge t_{1}}K_{0}ds$
$\leq V(\xi(0))+2K_{0}\tau,$
for all $k\geq k_{0}$. This also means that
$\mathbb{P}(\eta_{k}\leq 2\tau)\leq\frac{V(\xi(0))+2K_{0}\tau}{V(k)\wedge V(1/k)}.$
As $k\rightarrow\infty$, we get $\eta_{\infty}>2\tau$ almost surely. Meanwhile, for $t_{1}\in[0,T]$, we also derive from the Itô formula that
$\mathbb{E}(V(x(\eta_{k}\wedge t_{1})))$ $\leq V(\xi(0))+\mathbb{E}\int_{0}^{\eta_{k}\wedge t_{1}}K_{0}ds$
$\leq V(\xi(0))+K_{0}T,$
for all $k\geq k_{0}$. This also means that
$\mathbb{P}(\eta_{k}\leq T)\leq\frac{V(\xi(0))+K_{0}T}{V(k)\wedge V(1/k)}.$ (17)
As $k\rightarrow\infty$, we have $\eta_{\infty}>T$ almost surely. Repeating this procedure for $t_{1}\in[0,\infty)$, it is easy to see that $\mathbb{P}(\eta_{\infty}\leq\infty)\rightarrow 0$ as $k\rightarrow\infty$. This implies that $\eta_{\infty}=\infty$ almost surely and hence we must have $\eta_{e}=\infty$ almost surely. Therefore, the solution remains strictly positive and finite almost surely on the interval $[0,T]$, that is,
$\mathbb{P}\big(0<x(t)<\infty\text{ for all }t\in[0,T]\big)=1.$
3.2 Boundedness
We also present the following useful result that is required to establish uniform boundedness of the true solution of SDDE (6).
Lemma 3.2.
Let Assumption 2.1 hold and $\beta\in(0,1)$. Then for any initial value $\xi(0)$, the solution $x(t)$ of SDDE (6) fulfils
$\mathbb{E}\Big[x(t)^{\beta}-1-\beta\log(x(t))\Big]\leq\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}$
for all $t\geq 0$ and
$\underset{t\rightarrow\infty}{\limsup}\mathbb{E}\Big[x(t)^{\beta}-1-\beta\log(x(t))\Big]\leq\bar{K}_{0}$
where $\bar{K}_{0}$ is a positive constant that does not depend on the initial value $\xi(0)$.
Proof.
We define $V_{1}\in C^{2,1}(\mathbb{R}_{+}\times\mathbb{R}_{+};\mathbb{R}_{+})$ by $V_{1}(x,t)=e^{t}V(x)$, where $V(x)$ is the same as (15). We compute from the diffusion operator in (8) that
$LV_{1}(x,y,t)$ $=e^{t}\Big[x^{\beta}-1-\beta\log(x)+\alpha\mu\beta x^{\beta-1}-\alpha\mu\beta x^{-1}-\alpha\beta x^{\gamma+\beta-1}+\alpha\beta x^{\gamma-1}$
$+\frac{\sigma^{2}}{2}\beta(\beta-1)x^{2\theta+\beta-2}y^{2r}+\frac{\sigma^{2}}{2}\beta x^{2\theta-2}y^{2r}\Big]$
$=e^{t}\Big[x^{\beta}-1-\beta\log(x)\Big]+e^{t}\Big[\alpha\mu\beta x^{\beta-1}-\alpha\mu\beta x^{-1}-\alpha\beta x^{\gamma+\beta-1}+\alpha\beta x^{\gamma-1}$
$+\frac{\sigma^{2}}{2}\beta(\beta-1)x^{2\theta+\beta-2}y^{2r}+\frac{\sigma^{2}}{2}\beta x^{2\theta-2}y^{2r}\Big].$
Hence, by (16), there exists a constant such that
$LV_{1}(x,y,t)$ $\leq e^{t}(V(x)+LV(x,y))$
$\leq e^{t}\bar{K}_{0}.$
Using the same stopping time as defined in (14), we derive from the Itô formula that
$\mathbb{E}V_{1}(x(t\wedge\tau_{k}),t\wedge\tau_{k})\leq V_{1}(\xi(0),0)+\int_{0}^{t\wedge\tau_{k}}LV_{1}(x(s),x(t-\tau),s)ds.$
This implies that
$e^{t\wedge\tau_{k}}V(x(t\wedge\tau_{k}))\leq V(\xi(0))+\int_{0}^{t\wedge\tau_{k}}e^{s}(V(x(s))+LV(x(s),x(t-\tau)))ds.$
By applying the Fatou lemma and setting $k\rightarrow\infty$, we now have
$e^{t}V(x(t))\leq V(\xi(0))+e^{t}\bar{K}_{0}.$
This also implies that
$x(t)^{\beta}-1-\beta\log(x(t))\leq\frac{V(\xi(0))}{e^{t}}+\bar{K}_{0},$
which gives both assertions as required. ∎
The following result reveals that the true solution of SDDE (6) will stay in a compact support with large probability.
Theorem 3.3.
Let Assumption 2.1 hold and $\beta\in(0,1)$. Then for any initial value $\xi(0)$ and $k>k_{0}$, there exists a constant $\bar{K}_{0}$ such that
$\mathbb{P}(1/k<x(t)<k)\geq 1-\epsilon$
for all $t\geq 0$,
$\epsilon=\Big[\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}\Big]\Big[\frac{1}{(1/k)^{\beta}-1+\beta\log(k)}+\frac{1}{k^{\beta}-1-\beta\log(k)}\Big].$
Proof.
We compute from Lemma 3.2 that for any $t\geq 0$, we have
$\mathbb{P}(x(t)$ $\leq 1/k)\leq\mathbb{E}\Big[1_{\{x(t)\leq 1/k\}}\frac{x(t)^{\beta}-1-\beta\log(x(t))}{(1/k)^{\beta}-1+\beta\log(k)}\Big]$
$\leq\frac{\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}}{(1/k)^{\beta}-1+\beta\log(k)}.$
Similarly, we also obtain from Lemma 3.2 that for any $t\geq 0$
$\mathbb{P}(x(t)$ $\geq k)\leq\mathbb{E}\Big[1_{\{x(t)\geq k\}}\frac{x(t)^{\beta}-1-\beta\log(x(t))}{k^{\beta}-1-\beta\log(k)}\Big]$
$\leq\frac{\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}}{k^{\beta}-1-\beta\log(k)}.$
This implies that,
$\mathbb{P}(1/k<x(t)<k)\geq 1-\Big[\xi(0)^{\beta}-1-\beta\log(\xi(0))+\bar{K}_{0}\Big]\Big[\frac{1}{(1/k)^{\beta}-1+\beta\log(k)}+\frac{1}{k^{\beta}-1-\beta\log(k)}\Big],$
as required. ∎
4 Numerical method
In this section, we develop truncated EM techniques to estimate the solution of SDDE (6). Moreover, we establish some properties of the numerical solutions.
4.1 The truncated EM method
Before we construct the numerical method, we need to extend the domain of SDDE (6) from $\mathbb{R}_{+}$ to $\mathbb{R}$. We should mention that this extension does not affect previous results in anyway. To define the truncated EM method, we choose a strictly increasing continuous function $z:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that $z(u)\rightarrow\infty$ as $u\rightarrow\infty$ and
$\sup_{|x|\vee|y|\leq u}\Big(|f(x)|\vee g(x,y)\Big)\leq z(u),$ (18)
for all $u\geq 0$. Denote by $z^{-1}$ the inverse function of $z$ and we see that $z^{-1}$ is strictly increasing continuous function from $z(0),\infty)$ to $\mathbb{R}_{+}$. We also choose a number $\Delta^{*}\in(0,1]$ and a strictly decreasing function $\psi:(0,\Delta^{*}]\rightarrow\mathbb{R}_{+}$ such that
$\quad\psi(\Delta^{*})\geq z(1),\lim_{\Delta\rightarrow 0}\psi(\Delta)=\infty\text{ and }\Delta^{1/4}\psi(\Delta)\leq 1,\quad\forall\Delta\in(0,1).$ (19)
For a given step size $\Delta\in(0,\Delta^{*})$, we then define the truncated functions by
$f_{\Delta}(x)$ $=\begin{cases}f\Big(x\wedge z^{-1}(\psi(\Delta))\Big),&\mbox{if $x\geq 0$ }\\ \alpha\mu,&\mbox{if $x<0$},\end{cases}$
$g_{\Delta}(x,y)$ $=\begin{cases}g\Big(x\wedge z^{-1}(\psi(\Delta)),y\wedge z^{-1}(\psi(\Delta))\Big),&\mbox{if $x,y\geq 0$ }\\ 0,&\mbox{if $x,y<0$},\end{cases}$
for all $x,y\in\mathbb{R}$. So for $x,y\in[0,z^{-1}(\psi(\Delta))]$, we observe that
$|f_{\Delta}(x)|\vee g_{\Delta}(x,y)\leq z(z^{-1}(\psi(\Delta)))=\psi(\Delta),$ (20)
for all $x,y\in\mathbb{R}$. That is, $f_{\Delta}$ and $g_{\Delta}$ are bounded by $\psi(\Delta)$ although $f$ and $g$ are unbounded. From now on, we let $T>0$ be arbitrarily fixed. We also let the step size $\Delta\in(0,\Delta^{*}]$ be a fraction of $\tau$, that is, $\Delta=\frac{\tau}{M}$ for some integer $M>\tau$. We construct the discrete-time truncated EM approximation of SDDE (6) by defining $t_{k}=k\Delta$ for $M\leq k\leq\infty$, setting $X_{\Delta}(t_{k})=\xi(t_{k})$ for $M\leq k\leq 0$ and computing
$X_{\Delta}(t_{k+1})=X_{\Delta}(t_{k})+f_{\Delta}(X_{\Delta}(t_{k}))\Delta+g_{\Delta}(X_{\Delta}(t_{k}),X_{\Delta}(t_{k-M}))\Delta B_{k}$ (21)
for $k\geq 0$, where $\Delta B_{k}=B(t_{k+1})-B(t_{k})$ is an increment of the Brownian motion. For $t\in[-\tau,\infty)$, we define the continuous-time truncated EM step process by
$\bar{x}_{\Delta}(t)=\sum_{k=-M}^{\infty}X_{\Delta}(t_{k})1_{[t_{k},t_{k+1})}(t)$ (22)
where $1_{[t_{k},t_{k+1})}$ is the indicator function on $t_{k},t_{k+1})$. The continuous-time continuous truncated EM process is defined by setting $x_{\Delta}(t)=\xi(t)$ for $t\in[-\tau,0]$ while for $t\geq 0$, we get
$x_{\Delta}(t)=\xi(0)+\int_{0}^{t}f_{\Delta}(\bar{x}_{\Delta}(s))ds+\int_{0}^{t}g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))dB(s).$ (23)
We observe that $x_{\Delta}(t)$ is an Itô process on $t\geq 0$ satisfying Itô differential
$dx_{\Delta}(t)=f_{\Delta}(\bar{x}_{\Delta}(t))dt+g_{\Delta}(\bar{x}_{\Delta}(t),\bar{x}_{\Delta}(t-\tau))dB(t).$ (24)
It is important to note that $x_{\Delta}(t_{k})=\bar{x}_{\Delta}(t_{k})=X_{\Delta}(t_{k})$ for $k\geq-M$.
4.2 Properties of numerical solution
The following lemma shows that the discrete-time process $\bar{x}_{\Delta}(t)$ and the continuous-time process $x_{\Delta}(t)$ are close to each other in the strong sense.
Lemma 4.1.
For any fixed $\Delta\in(0,\Delta^{*}]$ and $p\geq 2$, we have
$\mathbb{E}|x_{\Delta}(t)-\bar{x}_{\Delta}(t)|^{p}\leq c_{p}\Delta^{p/2}(\psi(\Delta))^{p},$ (25)
for all $t\geq 0$, where $c_{p}$ is a generic constant that is dependent only on $p$.
Proof.
Fix any $\Delta\in(0,\Delta^{*}]$ and $t\geq 0$. Then there is a unique integer $k\geq 0$ such that $t_{k}\leq t\leq t_{k+1}$. By elementary inequality and (20), we have
$\mathbb{E}|x_{\Delta}(t)-\bar{x}_{\Delta}(t)|^{p}=\mathbb{E}|x_{\Delta}(t)-\bar{x}_{\Delta}(t_{k})|^{p}$
$\leq C_{p}\Big(\mathbb{E}\big|\int_{t_{k}}^{t}f_{\Delta}(\bar{x}_{\Delta}(s))ds\big|^{p}+\mathbb{E}\big|\int_{t_{k}}^{t}g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))dB(s)\big|^{p}\Big)$
$\leq C_{p}\Big(\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t}|f_{\Delta}(\bar{x}_{\Delta}(s))|^{p}ds+\Delta^{(p-2)/2}\mathbb{E}\int_{t_{k}}^{t}|g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))|^{p}ds\Big)$
$\leq C_{p}\Big(\Delta^{p-1}(\psi(\Delta))^{p}\Delta+\Delta^{(p-2)/2}(\psi(\Delta))^{p}\Delta\Big)$
$\leq C_{p}\Big(\Delta^{p}(\psi(\Delta))^{p}+\Delta^{p/2}(\psi(\Delta))^{p}\Big)\leq c_{p}\Delta^{p/2}(\psi(\Delta))^{p},$
where $c_{p}=C_{p}\vee 1$ and from (19), we obtain $\Delta^{p/2}(\psi(\Delta))^{p}\leq\Delta^{p/4}$. ∎
The following lemma reveals the probability that the truncated EM solutions do not explode in finite time.
Lemma 4.2.
Let Assumptions 2.1 and 2.1 hold and $T>0$ be fixed. For any sufficiently large integer $k>0$, define the stopping time by
$\eta_{\Delta}=\inf\{t\in[0,T]:x_{\Delta}(t)\notin[1/k,k]\}.$ (26)
Then for any fixed $\Delta\in(0,\Delta^{*}]$, we have
$\mathbb{P}(\eta_{\Delta}\leq T)\leq\frac{V(\xi(0))+K_{1}T+K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)},$ (27)
where $K_{1}$, $K_{2}$ and $K_{3}$ are generic constants and $V$ is defined in (15).
Proof.
For $t_{1}\in[0,T]$, we apply the Itô formula to (24) to compute
$\mathbb{E}(V(x_{\Delta}(t\wedge\eta_{\Delta})))-V(\xi(0))$
$=\mathbb{E}\int_{0}^{t_{1}\wedge\eta_{\Delta}}\Big(V_{x}(x_{\Delta}(s))f_{\Delta}(\bar{x}_{\Delta}(s))+\frac{1}{2}V_{xx}(x_{\Delta}(s))g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))^{2}\Big)ds$
$\leq\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}\Big(V_{x}(x_{\Delta}(s))f_{\Delta}(x_{\Delta}(s))+\frac{1}{2}V_{xx}(x_{\Delta}(s))g_{\Delta}(x_{\Delta}(s),x_{\Delta}(s-\tau))^{2}\Big)ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}V_{x}(x_{\Delta}(s))\Big(f_{\Delta}(\bar{x}_{\Delta}(s))-f_{\Delta}(x_{\Delta}(s))\Big)ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}\frac{1}{2}V_{xx}(x_{\Delta}(s))\Big(g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))^{2}-g_{\Delta}(x_{\Delta}(s),x_{\Delta}(s-\tau))^{2}\Big)ds.$
By recalling the definition of the truncated functions in (18), we note that
$f_{\Delta}(\cdot)=f(\cdot)\text{ and }g_{\Delta}(\cdot,\cdot)=g(\cdot,\cdot).$ (28)
Also, for $s\in[0,\eta_{\Delta}\wedge t_{1}]$ with $x_{\Delta}(s),\bar{x}_{\Delta}(s),x_{\Delta}(s-\tau),\bar{x}_{\Delta}(s-\tau)\in[1/k,k]$, we observe that
$g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))\vee g(x_{\Delta}(s),x_{\Delta}(s-\tau))\leq z(k).$ (29)
By Assumption 2.2, we have
$\mathbb{E}(V(x_{\Delta}(t\wedge\eta_{\Delta})))-V(\xi(0))\leq K_{1}T+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}V_{x}(x_{\Delta}(s))|f(\bar{x}_{\Delta}(s))-f(x_{\Delta}(s))|ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}\frac{1}{2}V_{xx}(x_{\Delta}(s))\Big(g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))^{2}-g(x_{\Delta}(s),x_{\Delta}(s-\tau))^{2}\Big)ds,$
where $LV(x_{\Delta}(s),x_{\Delta}(s-\tau))\leq K_{1}$ for $s\in[0,t\wedge\eta_{\Delta}]$. By an elementary inequality, we have
$\mathbb{E}(V(x_{\Delta}(t\wedge\eta_{\Delta})))-V(\xi(0))\leq K_{1}T+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}V_{x}(x_{\Delta}(s))|f(\bar{x}_{\Delta}(s))-f(x_{\Delta}(s))|ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}\frac{1}{2}V_{xx}(x_{\Delta}(s))\Big(|g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))-g(x_{\Delta}(s),x_{\Delta}(s-\tau))|$
$\times|g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))+g(x_{\Delta}(s),x_{\Delta}(s-\tau))|\Big)ds.$
By (20), (28), (29) and Lemma 2.3, we now have
$\mathbb{E}(V(x_{\Delta}(t\wedge\eta_{\Delta})))$ $\leq V(\xi(0))+K_{1}T+\mathbb{E}\int_{0}^{t_{1}\wedge\eta_{\Delta}}G_{k}V_{x}(x_{\Delta}(s))|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}(z(k))^{2}G_{k}V_{xx}(x_{\Delta}(s))|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$+\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}(z(k))^{2}G_{k}V_{xx}(x_{\Delta}(s))|\bar{x}_{\Delta}(s-\tau)-x_{\Delta}(s-\tau)|ds$
$\leq V(\xi(0))+K_{1}T+K_{2}\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$+K_{3}\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s-\tau)-x_{\Delta}(s-\tau)|ds,$
where
$K_{2}=\max_{1/k\leq x\leq k}\Big(G_{k}V_{x}(x)+(z(k))^{2}G_{k}V_{xx}(x)\Big)$
and
$K_{3}=\max_{1/k\leq x\leq k}\Big((z(k))^{2}G_{k}V_{xx}(x)\Big).$
So by Assumption 2.1 and Lemma 4.1, we get
$\mathbb{E}(V(x_{\Delta}(t\wedge\eta_{\Delta})))$ $\leq V(\xi(0))+K_{1}T+K_{2}\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$+K_{3}\mathbb{E}\int_{-\tau}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$\leq V(\xi(0))+K_{1}T+K_{2}\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$+K_{3}\mathbb{E}\int_{-\tau}^{0}|\xi([s/\Delta]\Delta)-\xi(s)|ds+K_{3}\mathbb{E}\int_{0}^{\eta_{\Delta}\wedge t_{1}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|ds$
$\leq V(\xi(0))+K_{1}T+K_{3}\int_{-\tau}^{0}\mathbb{E}|\xi([s/\Delta]\Delta)-\xi(s)|ds$
$+(K_{2}+K_{3})\int_{0}^{T}(\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p})^{1/p}ds$
$\leq V(\xi(0))+K_{1}T+K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T.$
This implies that
$\mathbb{P}(\eta_{\Delta}\leq T)\leq\frac{V(\xi(0))+K_{1}T+K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)},$
as required. ∎
5 Convergence analysis
In this section, we study the finite-time convergence of the truncated EM solutions to the true solution of SDDE (6). Further, we show that the truncated EM solutions converge to the true solution of SDDE (6) in probability. We perform simulation examples to support the findings and justify the convergence result for valuing some option contracts.
5.1 Finite-time error bound
The following lemma shows that the truncated EM solutions converge to the true solution of SDDE (6) in finite time.
Lemma 5.1.
Let Assumptions 2.1 and 2.2 hold. Then for any $p\geq 2$, fixed $T>0$, sufficiently large $k>0$ and $\Delta\in(0,\Delta^{*}]$, we have
$\mathbb{E}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)\leq K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p}),$ (30)
where $K_{8}$ is a generic constant that depends on $k$ but is independent of $\Delta$ and $\eta^{\Delta}_{k}=\eta_{k}\wedge\eta_{\Delta}$, where $\eta_{k}$ and $\eta_{\Delta}$ are defined in (14) and (26) respectively. Consequently, we have
$\lim_{\Delta\rightarrow 0}\mathbb{E}\Big(\sup_{0\leq t\leq T}|\bar{x}_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)=0.$ (31)
Proof.
It follows from (7) and (24) that for $t\in[0,t_{1}]$, we have
$\mathbb{E}\Big(\sup_{0\leq t\leq t_{1}}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)\leq K_{4}+K_{5},$
where
$K_{4}$ $=2^{p-1}\Big(\mathbb{E}\Big|\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}(f_{\Delta}(\bar{x}_{\Delta}(s))-f(x(s)))ds\Big|^{p}\Big)$
$K_{5}$ $=2^{p-1}\Big(\mathbb{E}(\sup_{0\leq t\leq t_{1}}\Big|\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k})}(g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))-g(x(s),x(s-\tau)))dB(s)\Big|^{p})\Big).$
By the Hölder inequality and (28), we compute
$K_{4}$ $\leq 2^{p-1}T^{p-1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|f_{\Delta}(\bar{x}_{\Delta}(s))-f(x(s))|^{p}ds\Big)$
$\leq 2^{p-1}T^{p-1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|f(\bar{x}_{\Delta}(s))-f(x(s))|^{p}ds\Big)$
$\leq 2^{p-1}T^{p-1}G_{k}^{p}\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|\bar{x}_{\Delta}(s)-x(s)|^{p}ds.$
Moreover, by the elementary inequality, we now have
$K_{4}$ $\leq c_{0}\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds+c_{0}\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|x_{\Delta}(s)-x(s)|^{p}ds$
$\leq c_{0}\int_{0}^{T}\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds+c_{0}\int_{0}^{t_{1}}\mathbb{E}\Big(\sup_{0\leq t\leq s}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)ds,$ (32)
where $c_{0}=2^{2(p-1)}T^{p-1}G_{k}^{p}$. By the Burkholder-Davis-Gundy inequality, Lemma 2.3 and (28), we also have
$K_{5}$ $\leq 2^{p-1}T^{\frac{p-2}{2}}\bar{c}_{p}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|g_{\Delta}(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))-g(x(s),x(s-\tau))|^{p}ds\Big)$
$\leq 2^{p-1}T^{\frac{p-2}{2}}\bar{c}_{p}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))-g(x(s),x(s-\tau))|^{p}ds\Big)$
where $\bar{c}_{p}$ is a positive constant. By the elementary inequality and Lemma 2.3, we get
$K_{5}$ $\leq 2^{2(p-1)}T^{\frac{p-2}{2}}\bar{c}_{p}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|g(\bar{x}_{\Delta}(s),\bar{x}_{\Delta}(s-\tau))-g(x_{\Delta}(s),x_{\Delta}(s-\tau))|^{p}ds\Big)$
$+2^{2(p-1)}T^{\frac{p-2}{2}}\bar{c}_{p}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|g(x_{\Delta}(s),x_{\Delta}(s-\tau))-g(x(s),x(s-\tau))|^{p}ds\Big)$
$\leq c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds\Big)+c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|\bar{x}_{\Delta}(s-\tau)-x_{\Delta}(s-\tau)|^{p}ds\Big)$
$+c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|x_{\Delta}(s)-x(s)|^{p}ds\Big)+c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|x_{\Delta}(s-\tau)-x(s-\tau)|^{p}ds\Big),$
where $c_{1}=2^{3(p-1)}T^{\frac{p-2}{2}}\bar{c}_{p}G^{p}_{k}$. This also means that
$K_{5}$ $\leq c_{1}\int_{0}^{T}\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds+c_{1}\int_{-\tau}^{0}\mathbb{E}|\xi([s/\Delta]\Delta)-\xi(s)|^{p}ds$
$+c_{1}\int_{0}^{T}\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds+c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|x_{\Delta}(s)-x(s)|^{p}ds\Big)$
$+c_{1}\int_{-\tau}^{0}\mathbb{E}|\xi([s/\Delta]\Delta)-\xi(s)|ds+c_{1}\Big(\mathbb{E}\int_{0}^{t_{1}\wedge\eta^{\Delta}_{k}}|x_{\Delta}(s)-x(s)|^{p}ds\Big)$
$\leq 2c_{1}\int_{-\tau}^{0}\mathbb{E}|\xi([s/\Delta]\Delta)-\xi(s)|^{p}ds+2c_{1}\int_{0}^{T}\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds$
$+2c_{1}\int_{0}^{t_{1}}\mathbb{E}\Big(\sup_{0\leq t\leq s}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)ds.$ (33)
By combining $K_{4}$ and $K_{5}$, that is (5.1) and (5.1), we now have
$\mathbb{E}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)\leq 2c_{1}\int_{-\tau}^{0}\mathbb{E}|\xi([s/\Delta]\Delta)-\xi(s)|^{p}ds$
$+(c_{0}+2c_{1})\int_{0}^{T}\mathbb{E}|\bar{x}_{\Delta}(s)-x_{\Delta}(s)|^{p}ds+(c_{0}+2c_{1})\int_{0}^{t_{1}}\mathbb{E}\Big(\sup_{0\leq t\leq s}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)ds.$
So by Assumption 2.1, Lemma 4.1 and the Gronwall inequality, we obtain the required assertion as
$\mathbb{E}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)$
$\leq K_{6}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})+K_{7}\int_{0}^{t_{1}}\mathbb{E}\Big(\sup_{0\leq t\leq s}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big)ds$
$\leq K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p}),$
where $K_{6}=2c_{1}D+c_{p}(c_{0}+2c_{1})$, $K_{7}=c_{0}+2c_{1}$ and $K_{8}=K_{6}e^{K_{7}}$. Moreover, by Lemma 4.1, we also get (31) by letting $\Delta\rightarrow 0$. ∎
5.2 Convergence in probability
The following theorem shows that the truncated EM solutions converge to the true solution of SDDE (7) in probability.
Theorem 5.2.
Let $x(t)$ and $x_{\Delta}(t)$ be the true solution and the truncated EM solution of (7) and (24) respectively. Then for any fixed $T>0$, $\Delta\in(0,\Delta^{*}]$ and $p\geq 2$, we have
$\lim_{\Delta\rightarrow 0}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\Big)=0\text{ in probability}.$ (34)
and consequently
$\lim_{\Delta\rightarrow 0}\Big(\sup_{0\leq t\leq T}|\bar{x}_{\Delta}(t)-x(t)|^{p}\Big)=0\text{ in probability},$ (35)
where $\bar{x}_{\Delta}(t)$ is defined in (22).
Proof.
For arbitrarily small constants $\epsilon$ and $\lambda$, set
$\bar{\Omega}=\Big\{\omega:\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\geq\lambda\Big\}.$
Then
$\lambda\mathbb{P}(\bar{\Omega}\cap(\eta^{\Delta}_{k}\geq T))$ $=\lambda\mathbb{E}\Big(1_{(\eta^{\Delta}_{k}\geq T)}1_{\bar{\Omega}}\Big)$
$\leq\mathbb{E}\Big(1_{(\eta^{\Delta}_{k}\geq T)}\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\Big)$
$\leq\mathbb{E}\Big(\sup_{0\leq t\leq T\wedge\eta^{\Delta}_{k}}|x_{\Delta}(t)-x(t)|^{p}\Big)$
$\leq\mathbb{E}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t\wedge\eta^{\Delta}_{k})-x(t\wedge\eta^{\Delta}_{k})|^{p}\Big).$
By Lemma 5.1, we get
$\mathbb{P}(\bar{\Omega}\cap(\eta^{\Delta}_{k}\geq T))\leq\frac{K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})}{\lambda}.$ (36)
Furthermore, we compute
$\mathbb{P}(\bar{\Omega})$ $\leq\mathbb{P}(\bar{\Omega}\cap(\eta^{\Delta}_{k}\geq T))+\mathbb{P}(\eta^{\Delta}_{k}\leq T)$
$\leq\mathbb{P}(\bar{\Omega}\cap(\eta^{\Delta}_{k}\geq T))+\mathbb{P}(\eta_{k}\leq T)+\mathbb{P}(\eta_{\Delta}\leq T).$ (37)
So, by substituting (17), (27) and (36) into (5.2), we have
$\mathbb{P}(\bar{\Omega})$ $\leq\frac{V(\xi(0))+K_{0}T}{V(k)\wedge V(1/k)}+\frac{K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})}{\lambda}$
$+\frac{V(\xi(0))+K_{1}T+K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)}.$ (38)
Therefore, we can select $k$ sufficiently large such that
$\frac{2V(\xi(0))+K_{0}T+K_{1}T}{V(k)\wedge V(1/k)}\leq\frac{\epsilon}{2}$ (39)
and select $\Delta$ so small such that
$\frac{K_{3}D\Delta^{\ell}+c_{p}(K_{2}+K_{3})\Delta^{1/2}\psi(\Delta)T}{V(1/k)\wedge V(k)}+\frac{K_{8}(\Delta^{\ell}\vee\Delta^{p/2}(\psi(\Delta))^{p})}{\lambda}\leq\frac{\epsilon}{2}.$ (40)
So by combining (39) and (40), we now have
$\mathbb{P}\Big(\sup_{0\leq t\leq T}|x_{\Delta}(t)-x(t)|^{p}\geq\lambda\Big)\leq\epsilon,$ (41)
as desired. However, by Lemma 4.1, we also obtain (35) by setting $\Delta\rightarrow 0$. ∎
5.3 Numerical simulation
In this illustrative simulation example, we compare the performance of the truncated EM method (TEM) constructed for SDDE (7) with the backward EM method (BEM). We should clarify that, to the best of our knowledge, there exist no relevant literature for the numerical treatment of SDDE (7) based on the backward EM method. This illustration is just for the purpose of comparison. For the sake of simplicity, let us consider the following form of SDDE (7) given by
$dx(t)=4(2-x(t)^{2})dt+0.5x(t-2)^{2/3}x(t)^{3/5}dB(t),$ (42)
with the initial data $\xi(0)=0.2$, where $\tau=2$, $\gamma=2$, $r=2/3$ and $\theta=3/5$. Clearly, we see that Assumption 2.1 is satisfied. Moreover, we note that
$\sup_{|x|\vee|y|\leq u}\Big(|f(x)|\vee g(x,y)\Big)\leq 6.5u^{2}$
for all $u\geq 1$. This means that we have $z(u)=6.5u^{2}$ with inverse $z^{-1}(u)=(u/6.5)^{1/2}$. If we choose $\psi(\Delta)=\Delta^{-2/3}$, then $z^{-1}(\psi(\Delta))=(\Delta^{-2/3}/6.5)^{1/2}$. Table 1 and Figure 1 show empirical distributions of the TEM and BEM solutions using $\Delta=10^{-2}$. The plot of convergence of the TEM and BEM solutions is depicted in Figure 2 using the same step size. From the plots, we see that the solution paths of both methods are almost the same. By using the step sizes $10^{-2}$, $10^{-3}$, $10^{-4}$ and $10^{-5}$, we have the errors between the TEM and BEM solutions with a reference line of order 1 in Figure 3.
Numerical method min mean sd kurt skew max
TEM 0.0000 1.3720 0.1880 14.5202 -1.7707 1.9380
BEM 0.0000 1.3720 0.1825 12.9464 -2.0372 1.9010
Table 1: Empirical distribution of the TEM and BEM solutions
Figure 1: Plot of the empirical distribution of the TEM and BEM solutions
Figure 2: Convergence of the TEM and BEM solutions
Figure 3: Errors between the TEM and BEM solutions
5.4 Option valuation
We justify the convergence result for valuation of a bond and a lookback put option via efficient use of the Monte Carlo method.
Lemma 5.3.
Let $x(t)$ and $\bar{x}_{\Delta}(t)$ be the true solution and the truncated EM step solution of (7) and (24) respectively. If a bond price $B(T)$ at maturity time $T$ is given by
$B(T)=\mathbb{E}\Big[\exp\Big(-\int_{0}^{T}x(t)dt\Big)\Big],$ (43)
then an approximation of $B(T)$ is computed by
$B_{\Delta}(T)=\mathbb{E}\Big[\exp\Big(-\int_{0}^{T}\bar{x}_{\Delta}(t)dt\Big)\Big].$
So, by the Theorem 5.2, we get
$\lim_{\Delta\rightarrow 0}|B_{\Delta}(T)-B(T)|=0.$
Proof.
Let $\epsilon,\delta\in(0,1)$ be arbitrarily small. It is sufficient to prove that
$\mathbb{P}\Big[\Big|\exp\Big(-\int_{0}^{T}x(t)dt\Big)-\exp\Big(-\int_{0}^{T}\bar{x}_{\Delta}(t)dt\Big)\Big|\geq\delta\Big]<\epsilon.$
Using the inequality $\exp(-|x|)-\exp(-|y|)\leq|x-y|$, we have
$\Big|\exp\Big(-\int_{0}^{T}x(t)dt\Big)-\exp\Big(-\int_{0}^{T}\bar{x}_{\Delta}(t)dt\Big)\Big|$ $\leq\Big|\int_{0}^{T}[x(t)-\bar{x}_{\Delta}(t)]dt\Big|$
$\leq T\sup_{0\leq t\leq T}|x(t)-\bar{x}_{\Delta}(t)|.$
By applying Theorem 5.2, we obtain the desired assertion. ∎
Lemma 5.4.
Let $x(t)$ and $\bar{x}_{\Delta}(t)$ be the true solution and the truncated EM step solution of (7) and (24) respectively. If the expected payoff of the fixed strike lookback put option with strike $K$ is defined by
$P=\mathbb{E}\big[(K-\underset{0\leq t\leq T}{\inf}x(t))^{+}\big],$
then the approximate expected payoff based on $\bar{x}_{\Delta}(t)$ is
$P_{\Delta}=\mathbb{E}\big[(K-\underset{0\leq t\leq T}{\inf}\bar{x}_{\Delta}(t))^{+}\big].$
So, by the Theorem 5.2, we have
$\underset{\Delta\rightarrow 0}{\lim}|P-P_{\Delta}|=0.$
Proof.
In other words, we need to prove that
$\lim_{\Delta\rightarrow 0}|(K-\inf_{0\leq t\leq T}x(t))^{+}-(K-\inf_{0\leq t\leq T}|\bar{x}_{\Delta}(t)|)^{+}|=0\quad\text{in probability}.$
This also means that the theorem holds as long as we can establish that for any small constants $\epsilon>0$ and $\delta\in(0,1)$
$\mathbb{P}(|(K-\inf_{0\leq t\leq T}x(t))^{+}-(K-\inf_{0\leq t\leq T}|\bar{x}_{\Delta}(t)|)^{+}|\geq\delta)<\epsilon$ (44)
holds for all sufficiently small $\Delta$. We observe that
$|(K-\inf_{0\leq t\leq T}x(t))^{+}-(K-\inf_{0\leq t\leq T}|\bar{x}_{\Delta}(t)|)^{+}|$ $\leq|\inf_{0\leq t\leq T}x(t)-\inf_{0\leq t\leq T}|\bar{x}_{\Delta}(t)||$
$\leq\sup_{0\leq t\leq T}|x(t)-|\bar{x}_{\Delta}(t)||$
$\leq\sup_{0\leq t\leq T}|x(t)-\bar{x}_{\Delta}(t)|.$ (45)
Then, it follows that
$\mathbb{P}(|(K-\inf_{0\leq t\leq T}x(t))^{+}-(K-\inf_{0\leq t\leq T}|\bar{x}_{\Delta}(t)|)^{+}|\geq\delta)\leq\mathbb{P}(\sup_{0\leq t\leq T}|x(t)-\bar{x}_{\Delta}(t)|\geq\delta).$ (46)
So, by Theorem 5.2, we now have
$\mathbb{P}(\sup_{0\leq t\leq T}|x(t)-\bar{x}_{\Delta}(t)|\geq\delta)<\epsilon$ (47)
for all sufficiently small $\Delta$. So by combining (46) and (47) gives us (44). ∎
Acknowledgement
The author would like to acknowledge the financial support from the Heilbronn Institute for Mathematical Research (HIMR) and the UKRI/EPSRC Additional Funding Programme for Mathematical Sciences. The author would also like to thank Prof Mao Xuerong for the discussion about this work.
Declarations
Conflict of interest: The author declares that he has no conflict of interest.